Solve this equation: (I am checking my homework to see if i got it right)
4[3-3(x+1)]+1=2(-31-x)2x+31
Here's my work/answer:
12-12x-12+1=-62-2x+2x+31
-12x+1=-50-2x+2x=31
-12x+1=-50-2x+2x+19
-12x=-50-2x+2x+18
-12x=-32
x= 2 2/3
: Solve this equation: (I am checking my homework to see if i got it right)
4[3-3(x+1)]+1=2(-31-x)2x+31
Here's my work/answer:
12-12x-12+1=-62-2x+2x+31
-12x+1=-50-2x+2x=31
-12x+1=-50-2x+2x+19
-12x=-50-2x+2x+18
-12x=-32
x= 2 2/3

Answer by Alan3354(1926)

4[3-3(x+1)]+1=2(-31-x)2x+31
Here's my work/answer:
12-12x-12+1=-62-2x+2x+31
-12x+1=-50-2x+2x=31
-12x+1=-50-2x+2x+19
-12x=-50-2x+2x+18
-12x=-32
x= 2 2/3
------------------
4[3-3(x+1)]+1=2(-31-x)2x+31
4[3-3x-3]+1 = -62-2x + 2x+31 (you're missing a plus sign above)
-12x + 1 = -31
-12x = -32
x = 8/3 = 2 2/3